OpenStudy (babyslapmafro):

Please help me simplify the following function. (sqrt(x)-2)/(x-4)

OpenStudy (anonymous):

\[\frac{ (\sqrt{x}-2) }{ (x-4) } * \frac{ (\sqrt{x}+2) }{ (\sqrt{x}+2) } \rightarrow \frac{ (x-4) }{ (x-4)(\sqrt{x}+2) } = \frac{ 1 }{ \sqrt{x}+2 } \]